### v1

# Advancing Structure Prediction of Biomolecular Interaction via Contact-Guided Sampling with HelixFold-S1

Abstract
The abstract serves both as a general introduction to the topic and as a brief, non-technical summary of the main results and their implications. Authors are advised to check the author instructions for the journal they are submitting to for word limits and if structural elements like subheadings, citations, or equations are permitted.
Abstract
Biomolecular structure prediction is essential to molecular biology, yet accurately predicting the structures of complexes remains challenging, especially when co-evolutionary signals are absent. While recent methods have improved prediction accuracy through extensive sampling, aimless sampling often provides diminishing returns due to limited conformational diversity. Here, we introduce HelixFold-S1, a contact-guided sampling strategy that improves structural accuracy. Rather than relying on indiscriminate sampling, HelixFold-S1 predicts contact probabilities between molecular entities and uses these predictions to prioritize sampling of likely binding sites and modes. This targeted approach generates a diverse set of structural candidates, enhancing the likelihood of identifying accurate conformations. We demonstrate that HelixFold-S1 consistently outperforms baseline sampling strategies across a range of biomolecular interactions, including protein-antibody, protein–protein, protein–ligand, protein-RNA, and protein–DNA interfaces. Furthermore, the predicted contact probabilities serve as a reliable indicator of structural difficulty, guiding the allocation of sampling resources. These results highlight the potential of targeted sampling strategies to advance the structural modeling of complex biomolecular interactions during inference.
keywords:
Biomolecular interaction, Structure Prediction, Test-time scaling, Sampling Strategy
Introduction
Biomolecular structure prediction is a cornerstone of computational biology, underpinning advances in drug discovery, protein engineering, and our broader understanding of molecular interactions. Recent breakthroughs by deep learning-based methods [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11], represented by AlphaFold [1, 2, 3] and RoseTTAFold series [4, 5], have transformed the field by exploiting co-evolutionary signals and leveraging deep neural networks to infer complex structural relationships.
Despite these advances, predicting the structures of biomolecular complexes remains challenging, particularly when co-evolutionary signals are weak or unavailable. This is especially true for antigen–antibody pairs, where antibody diversity and the lack of co-evolution with antigens hinder accurate interface prediction. These limitations are further compounded by the difficulty of achieving high accuracy from a single predicted structure. While the AlphaFold series has sought to improve precision through model ensembling (e.g., five or twenty-five ensembles), the limited number of generated structures constrains comprehensive exploration of potential binding sites and interaction modes, ultimately restricting improvements in accuracy. To address this, recent studies [12, 13, 14, 15] have explored large-scale sampling strategies aimed at increasing conformational diversity by generating thousands of candidate structures. For example, AFSample [12] and AFSample2 [13] apply dropout and random MSA column masking during inference to produce diverse conformations, while AlphaFold3 [3] demonstrates that extensive sampling can significantly enhance predictions of antigen–antibody interfaces. These approaches share conceptual similarities with test-time scaling in large language models (LLMs) [16, 17, 18, 19, 20, 21], where greater computational resources are allocated during inference to improve output quality. Nevertheless, the structural variability of samples produced by folding models remains relatively low, leading to diminishing gains from repeated sampling. Accordingly, increasing the number of predictions alone provides only limited benefits in terms of accuracy. Consequently, merely increasing the number of predictions provides marginal improvements in accuracy. This highlights the necessity for more intelligent, resource-efficient sampling strategies that can prioritize the exploration of structurally informative regions and achieve higher accuracy with reduced computational overhead.
To address the inefficiencies of traditional large-scale sampling, we propose HelixFold-S1, which introduces a contact-guided sampling strategy to more efficiently explore the conformational space and improve structural prediction accuracy. Rather than exploring the conformational space randomly or exhaustively, HelixFold-S1 estimates a probability distribution over potential intermolecular contacts, prioritizing regions with a high likelihood of interaction. These predicted contact regions are used as spatial constraints to guide the model’s structural predictions, directing computational resources toward areas that provide the most structural insight. The generated conformations are then evaluated and ranked based on the associated confidence scores. This targeted approach enhances the structural diversity of the predictions, expanding the search for plausible binding modes and increasing the likelihood of identifying accurate structures.
We evaluate HelixFold-S1 across a range of biomolecular interaction scenarios, including protein–protein, protein–antibody, protein–ligand, protein–RNA, and protein–DNA interfaces. Compared to traditional sampling strategies, the contact-guided sampling approach employed by HelixFold-S1 demonstrates a marked improvement in prediction accuracy as the sampling process is increased. This advantage is particularly evident in more challenging scenarios, such as protein–antibody interactions with limited co-evolutionary signals and protein–ligand interfaces involving ligands absent from the training set. Moreover, the contact probabilities predicted by HelixFold-S1 exhibit a clear correlation with the difficulty of structural prediction. Specifically, targets with lower predicted contact probabilities tend to yield lower prediction accuracy, indicating greater structural complexity, while higher contact probabilities correspond to improved accuracy. For targets with intermediate contact probabilities, HelixFold-S1 shows a pronounced potential for improvement with increased sampling. Additionally, the conformational ensembles generated by HelixFold-S1 are more diverse, encompassing a broader spectrum of potential contact regions. This increased diversity enhances the robustness of the structural predictions and facilitates the identification of near-native conformations. HF-S1 has been deployed on the PaddleHelix platform and is available for online use.
Results
Framework of HelixFold-S1
Figure 1: Overall framework of HelixFold-S1. a, Inference pipeline of HF-S1: a three-stage process for contact-guided structure prediction. Stage 1: a contact prediction module generates inter-chain contact probabilities. Stage 2: an algorithm samples diverse contact subsets. Stage 3: selected contacts guide structure generation and ranking via the contact conditioning module. b, Network architecture of HF-S1. Input sequences, MSA/template features, and optional contact constraints are passed to the Input Embedder, with contact constraints processed by the contact conditioning module. The Pairformer models these inputs, followed by a diffusion module predicting the 3D structure. A contact prediction module estimates inter-token contact probabilities. During training, the contact conditioning and prediction modules are activated mutually exclusively.
HelixFold-S1 (HF-S1) is a sampling-optimized variant of HelixFold3 (HF3) [9], an open-source biomolecular structure prediction model built with PaddlePaddle [22] (Fig. 1). HF3 reproduces the core capabilities of AlphaFold3 (AF3) [3], supporting a broad range of biomolecules, including proteins, nucleic acids, and small molecules.
To enhance inference efficiency and structural diversity, HF-S1 introduces two contact-centric modules: the Contact Prediction Module and the Contact Conditioning Module (Fig. 1b). A contact is defined as existing between two tokens if any pair of atoms from the corresponding tokens lies within 5Å. The Contact Prediction Module computes a contact probability matrix across all token pairs using the pair representations produced by the Pairformer, a specialized attention-based module designed to capture long-range pairwise token dependencies. The Contact Conditioning Module, incorporated into the Input Embedder, enables the model to learn the contact constraints and generate structures based on given contact constraints. These two modules are activated in a mutually exclusive manner during training, creating a multi-task framework that alternates between contact probability estimation and contact-conditioned structure generation. This allows the model to jointly learn to infer inter-chain interactions and effectively apply them during inference.
The inference pipeline of HF-S1 consists of three stages (Fig. 1a). Contact Prediction: HF-S1 first computes a contact probability matrix over all token pairs, estimating the likelihood that any two tokens are in spatial proximity (i.e., any atom pair within 5Å). However, only the entries corresponding to inter-chain token pairs are considered valid, as the model is designed to focus on contacts between different molecular chains. This filtered matrix defines a probabilistic landscape of inter-chain interactions and guides the subsequent sampling process. Contact Sampling: To efficiently explore the structural landscape, HF-S1 employs a greedy sampling strategy that selects contacts in descending order of their predicted probabilities. Importantly, to reduce sampling redundancy, once a conformation is predicted using a particular contact as a constraint, any additional contacts that are also satisfied within the resulting conformation will be excluded from subsequent sampling. We call this strategy as redundant contact pruning. Contact-Guided Structure Prediction and Ranking: For each sampled contact, the model switches to contact conditioning mode to generate a structure that satisfies the contact constraint. All resulting structures are then ranked based on model confidence scores, and the top-ranked prediction is selected as the final output.
Improved Structural Accuracy across Complex Types
Figure 2: Improved structural accuracy across complex types. a, Representative example illustrating HF-S1’s sampling strategy. The heatmap depicts predicted inter-chain contact probabilities between residues of protein A and protein B. Surface residues are colored according to their highest predicted contact probability with the partner chain. b, Structural accuracy of HF-S1 evaluated on benchmark complexes collected from the RCSB PDB between January 1, 2022 and December 31, 2024. Accuracy improves with increased sampling, and HF-S1 consistently achieves high performance across a range of interface types, including protein–antibody (n=221), protein–protein (n=199), protein–ligand (n=194), protein–RNA (n=94), and protein–DNA (n=252). c, Comparison of HF-S1 to various baselines on protein-antibody complexes released in year 2024 (n=95) with sampling number of 80. d, Accuracy on protein-antibody complexes of year 2024 (n=95) changes with the sampling number.
To systematically evaluate the performance of HF-S1 across a range of biologically relevant complex types, we constructed a test set comprising protein–antibody (n=221), protein–protein (n=199), protein–ligand (n=194), protein–RNA (n=94), and protein–DNA (n=252) complexes, collected between 2022.01.01 and 2024.12.31 from the RCSB PDB [23]. To minimize overlap with the training data, all test samples were selected to have low sequence identity to the training set. Sequences were clustered by similarity [24], and one representative per cluster was randomly chosen to ensure diversity and reduce redundancy. For protein–ligand complexes, any ligands that appeared in the training data were further excluded.
We benchmarked HF-S1 against HF3 [9], an open-source implementation of AF3 that achieves accuracy comparable to the original model. For both models, multiple structural conformations were sampled per target. Prediction performance was evaluated using two selection strategies: score-based selection, in which the top-ranked conformation was chosen based on the model’s predicted confidence scores, and oracle selection, in which the most accurate conformation was retrospectively selected based on the experimental structure. For protein–protein complexes (including protein–antibody), performance was assessed using average DockQ scores [25]. For protein–ligand complexes, accuracy was measured by the proportion of predictions with a pocket-aligned ligand root mean squared deviation (RMSD) below 2Å. For protein–nucleic acid complexes (including protein–RNA and protein–DNA), interface accuracy was evaluated using interface LDDT (iLDDT) [26].
We first illustrate the model’s sampling strategy, we present a representative example of a complex structure predicted by HF-S1 (Fig. 2a). The model begins by generating an inter-chain contact probability matrix—visualized as a heatmap—that quantifies the likelihood of contact between each token pair from protein A and protein B. To highlight potential interaction regions, the surface residues of each protein are colored based on the highest contact probability in their respective rows or columns. This visualization effectively identifies residues most likely to participate in inter-chain interactions. Using this contact map as guidance, HF-S1 samples multiple plausible structural conformations, capturing diverse binding orientations between the two proteins.
Across all complex types (Fig. 2b), both HF3 and HF-S1 exhibited improved accuracy as the number of sampled conformations increased, underscoring the significance of test-time scaling in biological systems. The widening performance gap under oracle selection between HF-S1 and HF3 with additional conformations highlights the advantages of the advanced contact-guided sampling strategies implemented in HF-S1, which effectively enhance conformational diversity and structural accuracy. Notably, HF-S1 demonstrated particularly substantial improvements for protein–antibody interfaces compared with general protein–protein interfaces. HF-S1 yields greater performance gains over HF3 on protein–antibody complexes than on general protein–protein interactions (Fig. 2c). This is likely due to the absence of co-evolutionary signals in antibody–antigen interactions, which results from the highly diverse and individually generated nature of antibodies and the lack of long-term evolutionary coupling between binding partners. For protein–ligand complexes, the intrinsic flexibility and diverse binding modes of ligands pose additional challenges for traditional methods. In these scenarios, HelixFold-S1’s enhanced sampling mitigates the absence of strong evolutionary constraints, leading to improved prediction accuracy. Furthermore, HF-S1 exhibited greater improvements under the oracle selection strategy than under score-based selection, indicating that the current scoring function may not fully capture the most accurate conformations. This suggests that further optimization of the score-based selection process could help better exploit the benefits of enhanced conformational sampling.
We further benchmark HF-S1 against several representative baselines on protein–antibody complexes (Fig. 2c and 2d), including two state-of-the-art AF3-like structure prediction models, Chai-1 [27] and Boltz-2 [28], as well as the advanced sampling algorithm AFsample [12]. Among them, Chai-1 and Boltz-2 reflect the current leading performance in biomolecular structure prediction, while AFsample represents a strong post-hoc sampling enhancement method. To assess the impact of sampling strategies independently of model backbone quality, we construct a hybrid baseline, HF3 w/ AFsample, by replacing HF3’s original sampling method with that of AFsample. This enables a direct comparison of sampling effectiveness across methods. To ensure fair evaluation and avoid data leakage—particularly considering that Boltz-2 was trained on PDB entries up to June 1, 2023—we curate a test subset of 95 protein–antibody complexes released after January 1, 2024. All models are evaluated under the same condition of 80 samples per target. Using DockQ ¿ 0.23 as the success criterion (Fig. 2d), HF3 and Boltz-2 achieve strong baseline accuracy, both outperforming Chai-1. However, their performance shows limited improvement as the number of samples increases, indicating that their native sampling strategies are less effective at exploring alternative conformations. Chai-1, despite being a leading AF3-like model, shows virtually no gain from additional sampling, suggesting minimal responsiveness to increased sample size. By contrast, HF3 w/ AFsample yields noticeable accuracy gains with more samples, confirming the advantage of more sophisticated sampling techniques. Most notably, HF-S1 exhibits significantly faster and more substantial improvements than all baselines, highlighting its superior ability to efficiently explore the structural landscape and refine predictions through sampling.
Diversity and Quality of the Sampled Conformations
Figure 3: Diversity and quality of the sampled conformations. a, Distribution of structural accuracy across all sampled conformations at each sampling step, averaged over multiple datasets. We report the highest, lowest, and 25th-percentile precision scores to characterize both the range and quality of sampled structures. b, Representative examples showing that HF-S1 produces more diverse and higher-quality conformations than HF3, as reflected in the broader distributions of confidence and precision.
HF-S1 outperforms HF3 primarily due to its contact-guided sampling strategy, which stands in contrast to the unguided (random) sampling employed by HF3. We hypothesize that the key advantage of HF-S1 lies in its ability to generate a broader and higher-quality distribution of sampled conformations. Specifically, increased sampling diversity may raise the likelihood of capturing high-precision structures during the inference process.
To evaluate this hypothesis, we analyzed the distribution of structural accuracy across all sampled conformations at each sampling step. For each target and sampling num $s$, we sorted the $s$ generated conformations by their precision scores and tracked three statistics: the highest precision, the lowest precision, and the 25th-percentile precision (i.e., the precision of the conformation ranked at the top 25%). The averaged results across multiple datasets are presented in Fig. 3a. This analysis reveals two key patterns. First, as the number of samples increases, both HF3 and HF-S1 exhibit a broader spread in structural precision—indicating that later-stage sampling leads to both better and worse structures. Second, and more importantly, HF-S1 consistently produces a wider precision range compared to HF3, confirming that the contact-guided approach enhances structural diversity. Furthermore, the 25th-percentile precision values for HF-S1 are consistently higher than those of HF3, indicating that HF-S1 tends to generate a larger fraction of high-quality conformations.
We further illustrate the differences between HF3 and HF-S1’s sampling strategies using two representative examples (Fig. 3b). The first example is the complex structure of fungal $\beta$-1,3-glucanosyltransferases (Gel4) and Nb4 nanobody, where the nanobody binds to a dissimilar CBM43 domain of Gel4 across fungal species (PDB ID: 8pe1) [29]. The second example is a protein-RNA heterodimer, illustrating the interaction between the MD-4HB protein and helix 44 RNA in the yeast ribosome (PDB ID: 7x34) [30]. For each case, we examined the distribution of model-predicted confidence scores and interface-level precision metrics across the sampled conformations. HF-S1 consistently generates a broader spectrum of structures, spanning a wide range of confidence levels and precision values—including a notable fraction of high-accuracy predictions. In contrast, HF3 tends to sample conformations within a narrower confidence and precision range, indicating more limited structural diversity and fewer high-quality candidates.
Predicted Contact Probability
Figure 4: Predicted Contact Probability. a, Accuracy of the contact probability predicted by HF-S1. b, Contact satisfaction rate, i.e., fraction of structures predicted by HF-S1 satisfying contact constraints. c, Correlation between the target-level contact probability (maximum value in the predicted contact probability matrix) and the accuracy of the predicted structure. d, Improvement from multiple sampling by HF-S1 across different target contact probability groups under oracle selection and score-based selection.
The contact probability matrix predicted by HF-S1 serves a dual role: it not only informs the structural sampling strategy but also provides insight into the intrinsic difficulty of the structure prediction task, as well as the potential benefits of additional sampling.
We first evaluate the accuracy of the contact probability matrix outputted by HF-S1. For each target, token pairs that form true contacts, as defined by experimental structures, are treated as positive examples, while all others are considered negative. Using this approach, we calculate the area under the precision–recall curve (AUPRC) between the predicted and ground-truth contact maps, reporting the average values across all targets. As a baseline, for each target, we generate a posterior contact probability matrix for HF3 based on the sampled conformations. The matrix element for each token pair is defined as the inverse of the minimum distance observed across the sampled conformations. In comparison to this baseline, the contact probabilities predicted by HF-S1 consistently show strong accuracy across a variety of molecular types (Fig. 4a). Among these categories, protein–antibody complexes exhibit the lowest AUPRC, further confirming the difficulty of this prediction scenario. Interestingly, despite the lack of coevolutionary signals in protein–ligand binding, protein–ligand complexes achieve the highest AUPRC. This may be attributed to the fact that binding sites in protein–ligand interactions are often structurally conserved and spatially well-defined. Many such interfaces exhibit clear geometric features—such as hydrophobic pockets and polar residue arrangements—that help the model accurately localize the binding site.
We next assessed whether the conformations predicted by HF-S1 satisfy the input contact constraints. Specifically, we evaluated the contact satisfaction rate, defined as the fraction of predicted structures in which the specified contacts are realized (Fig. 4b). Across most test cases, HF-S1 consistently adhered to the provided constraints, achieving satisfaction rates above 70% across various types of targets, demonstrating the model’s ability to effectively incorporate such priors into structure prediction.
It is of interest to examine whether predicted contact probabilities can serve as a proxy for the intrinsic difficulty of a target. To this end, we analyzed the relationship between the target-level contact probability—defined as the highest value in the predicted contact probability matrix for each target—and the prediction precision of HF-S1 under single sampling (Fig. 4c). Across all benchmark datasets, we observed a strong correlation: targets with lower contact probabilities tend to exhibit lower structural accuracy, whereas those with higher values are generally predicted more precisely. This suggests that contact probability estimates may reflect the inherent difficulty of the prediction task. When grouped by complex type, protein–protein targets typically show higher target-level contact probabilities, while protein–antibody complexes cluster in the lower range, indicating greater structural uncertainty. Protein–ligand complexes consistently exhibit high contact probabilities, suggesting that HF-S1 can often localize ligand binding sites with high confidence. In contrast, protein–RNA and protein–DNA complexes display a broader distribution, reflecting greater variability in prediction difficulty across these categories.
Finally, we examined whether target-level contact probabilities are associated with the degree of improvement achieved through multiple sampling. Targets were grouped based on their target-level predicted contact probability, and we analyzed the precision improvements of multi-sample predictions relative to single-sample predictions (Fig. 4d). We found that targets with intermediate contact probabilities achieved the greatest improvements, while those with low or high probabilities saw more modest gains under oracle selection. Notably, although low-probability targets did benefit from sampling, their improvements were generally smaller than those in the intermediate group. This trend is intuitive: for targets with high contact probabilities, accurate structures can often be recovered from a single sample, leaving limited room for further enhancement. For low-probability targets, the predicted contact maps are weak across the board, suggesting that a much larger number of samples may be needed to identify accurate structures. In contrast, intermediate cases offer partial yet informative contact signals, enabling the model to better explore the structural landscape and refine its predictions through sampling. As protein–ligand complexes consistently exhibit high target-level contact probabilities, they did not follow this trend. Besides, under score-based selection, the overall trend becomes less pronounced, largely due to limitations in the model’s confidence scoring. These results highlight the need to further improve the model’s confidence scoring mechanism.
Discussion
HelixFold-S1 exhibits strong performance across diverse complex structure prediction tasks by integrating contact-aware sampling strategies. Empirical evaluations demonstrate that increasing the number of sampled conformations systematically improves prediction accuracy, with the most notable gains observed in challenging cases such as antigen–antibody interactions. These results highlight test-time scaling as a practical and effective strategy for enhancing structural prediction quality. Moreover, the model’s predicted contact probabilities offer interpretable insights into plausible binding modes, contributing to both improved efficiency and a deeper understanding of protein interactions.
Despite these advances, several challenges remain. One key limitation is the accurate estimation of structural confidence. While the model has made progress in generating diverse conformations, current confidence scores are not always reliable in identifying the best candidate structures. Improving these scores—through more sophisticated scoring functions or data-driven refinements—could enhance model usability by streamlining candidate selection and increasing prediction reliability. Another challenge lies in sampling efficiency. Although HelixFold-S1 already achieves competitive performance in this regard, the number of conformations required for particularly difficult targets remains substantial. Incorporating prior structural knowledge to guide sampling could help narrow the conformational search space, thereby reducing computational overhead while maintaining or improving prediction accuracy.
In summary, HelixFold-S1 represents a meaningful step forward in complex structure prediction, offering robust capabilities for both academic research and industrial deployment. Addressing the remaining challenges will be key to unlocking its full potential for accurate, efficient, and reliable structural modeling across a broad range of biological systems.
Method
Model Architecture
HF-S1 builds upon the HF3 architecture and is designed to support two complementary tasks: inter-chain contact prediction and contact-conditioned structure prediction. To this end, HF-S1 introduces two additional components: the Contact Prediction Module and the Contact Conditioning Module, which extend the base architecture to enable contact-level reasoning and constraint-based structure generation, respectively.
The contact prediction task aims to estimate the inter-chain contact distribution of a given protein complex. Various input features—including sequence, multiple sequence alignment (MSA), and template information—are first encoded and then processed by a Pairformer module to generate single and pair representations. These pairwise representations are subsequently passed to a dedicated Contact Prediction Module, which includes a Pairformer Stack and performs a binary classification task on each element of the pairwise representation to output contact probabilities for each inter-chain token pair. A contact is defined as the presence of any atom pair between two tokens within 5Å in 3D space. The resulting pairwise contact probability matrix captures the inter-chain contact distribution of the complex and serves as an informative intermediate representation. By operating in the simplified space of contacts—rather than directly in the complex, high-dimensional structural space—the model achieves greater computational efficiency and facilitates the contact-conditioned structure prediction task.
The contact-conditioned structure prediction task introduces a Contact Conditioning Module to incorporate external contact constraints. These constraints are represented as a binary matrix $\{c_{ij}\}$, where $c_{ij}\in{0,1}$ indicates whether token pair $(i,j)$ is in contact (1 if any atom pair is within 5Å, and 0 otherwise). This matrix is projected through a linear layer and then fused into the pairwise activations within the Input Embedder of the model. During training, for each complex, 0–10 inter-chain contacts are randomly sampled from the contact set extracted from its ground-truth structure and provided as input. During inference, contact constraints are sampled from the contact probability matrix produced by the contact prediction task; the specific sampling strategy is described in a later section. The model learns to utilize the provided contact information to enhance structure prediction accuracy.
Training Regime
Parameters of the newly introduced Contact Prediction Module and Contact Conditioning Module are randomly initialized, while the remaining parts of the model inherit weights from the pretrained HF3. Fine-tuning is performed using the same training dataset as HelixFold3, which includes Protein Data Bank (PDB) [23] structures released before September 30, 2021, supplemented with self-distillation data to enhance generalization. The training follows a three-stage fine-tuning strategy: the first stage focuses on the contact-conditioned structure prediction task to improve complex structure accuracy with inter-chain contact constraints; the second stage adds the contact prediction task, jointly optimizing the model for both tasks; the third stage follows the same setting of the second stage but extends to larger crop size.
In the first stage, the model is trained exclusively on the contact-conditioned structure prediction task. For each training sample, inter-chain contacts are extracted from experimentally determined complex structures to form a ground-truth contact set $\mathcal{C}$. This set contains all inter-chain token pairs where at least one atom from each token lies within 5Å in three-dimensional space. Contact conditioning is applied with 70% probability: 1–10 token pairs are uniformly sampled from $\mathcal{C}$ and provided as binary contact constraints. In the remaining 30% of samples, no contact constraints are used, which helps maintain the model’s ability to predict structures without external guidance.
In the second stage, the contact prediction task is introduced, and the model is fine-tuned on both tasks simultaneously. The contact prediction task aims to estimate the probability that each inter-chain token pair is in atomic contact, serving as a basis for generating contact constraints during inference. During this task, the Contact Conditioning Module is not activated, and the model relies solely on encoded sequence, MSA, and template features to infer contact patterns. To supervise this task, a binary classification loss is applied over all inter-chain token pairs:
$\mathcal{L}_{\text{contact}}=\frac{1}{|\mathcal{P}|}\sum_{(i,j)\in\mathcal{P}}\text{cross\_entropy}(p_{ij}^{contact},y_{ij}^{contact}).$
Here, $\mathcal{P}$ denotes the set of all token pairs $(i,j)$ such that token $i$ and token $j$ belong to different chains. $p_{ij}^{contact}$ is the predicted probability of contact between tokens $i$ and $j$. $y_{ij}^{contact}=1$ if $(i,j)\in\mathcal{C}$ (i.e., the token pair is in contact), and $y_{ij}^{contact}=0$ otherwise. During training, half of the samples are used for the contact-conditioned structure prediction task, following the protocol established in the first stage, while the other half are dedicated to training the contact prediction task, which guides the model to estimate inter-chain contact probability distributions.
The loss function largely follows the original AF3/HF3 formulation, with an additional contact loss term introduced during fine-tuning:
$\mathcal{L}_{\text{loss}}=\alpha_{\text{confidence}}\mathcal{L}_{\text{confidence}}+\alpha_{\text{diffusion}}\mathcal{L}_{\text{diffusion}}+\alpha_{\text{distogram}}\mathcal{L}_{\text{distogram}}+\alpha_{\text{contact}}\mathcal{L}_{\text{contact}},$
with hyperparameters $\alpha_{\text{confidence}}=0.01$, $\alpha_{\text{diffusion}}=4$, and $\alpha_{\text{distogram}}=0.3$. The contact loss coefficient $\alpha_{\text{contact}}$ is set to 1 during training samples used for the contact prediction task and 0 during samples used for the contact-conditioned structure prediction task. The definitions of all other loss terms remain consistent with those in AF3.
All stages use the Adam optimizer [31] with parameters $\beta_{1}=0.9$, $\beta_{2}=0.95$, and $\epsilon=10^{-8}$, and a learning rate of $2\times 10^{-4}$. The mini-batch size is fixed at 128 for all stages. The first fine-tuning stage consists of 10,000 training steps with a crop size of 384. The second fine-tuning stage extends to 20,000 steps, also with a crop size of 384. The third stage continues training for an additional 3,000 steps with an increased crop size of 640.
Inference Regime
The inference process of HF-S1 (illustrated in Fig. 1a) consists of three stages: Contact Prediction, Contact Sampling, and Contact-Guided Structure Prediction and Ranking.
In the Contact Prediction stage, the contact prediction task is executed five times to reduce prediction variance, producing five contact probability matrices. These matrices are averaged element-wise to generate the final contact probability matrix, where each element $p_{ij}^{contact}$ represents the predicted contact probability between tokens $i$ and $j$. Only inter-chain contact probabilities are retained, with intra-chain contacts set to zero. For protein–antibody complexes, contact sampling is performed exclusively between the antigen chain and each antibody chain (heavy and light), excluding contacts between heavy and light chains.
In the Contact Sampling stage, inter-chain contacts are selected sequentially in descending order according to their predicted contact probabilities. Each selected contact is used as a binary constraint in the subsequent structure prediction step to generate diverse candidate structures. To improve sampling efficiency and avoid redundancy, two strategies are adopted: redundant contact pruning and enriched sampling of previously identified contact sets. We denote the sets of contacts extracted from previously predicted structures as $C_{1},C_{2},\ldots$, where each $C_{k}$ corresponds to the contacts obtained from the $k$-th predicted structure, following the same ground-truth extraction method described earlier. During sampling, redundant contact pruning excludes any candidate contact that overlaps with contacts already present in the union of all previously sampled sets $\bigcup_{i=1}^{k-1}C_{i}$. Here, overlapping means the candidate contact appears in any previously extracted contact set. This ensures that each newly sampled contact introduces novel constraints and helps maintain diversity among sampled structures. As sampling progresses, the predicted contact probabilities of remaining candidates naturally decrease. When these probabilities fall below a threshold (set as $0.2\cdot\max_{i,j}p_{ij}^{contact}$), the benefit of exploring new low-confidence contacts diminishes. At this point, instead of sampling new contacts, the algorithm enriches sampling by iterating through the existing contact sets $C_{1},C_{2},\ldots$ in order. Contacts are drawn from these sets cyclically to further exploit high-confidence information until the total sampling budget $S$ is reached.
In the Contact-Guided Structure Prediction and Ranking stage, each sampled contact is treated as a binary constraint and passed into the contact-conditioned structure prediction task to generate a candidate structure. A confidence score, named ranking_confidence, is computed for each structure, and the final prediction is selected as the one with the highest confidence among all candidates. Drawing inspiration from the AF3, we define the confidence score as a weighted average of the pTM and ipTM scores, with an additional penalty term for structural clashes. The score is computed as follows:
[…]
where pTM represents the predicted TM-Score for the full complex, indicating the confidence for overall structural accuracy. ipTM represents the interface predicted TM-Score for the full complex, focusing on the accuracy of interfacial interactions. has_clash is a binary term indicating the presence of obvious clashes between polymer chains in the predicted structure. Detailed definitions of pTM, ipTM, and has_clash can be found in the AF3 paper [3].
We adopt consistent inference settings across structure prediction tasks, including our method and the baselines Boltz‑2 and Chai‑1. Each prediction is refined using 10 recycling iterations and 200 diffusion steps, where the diffusion module is run once to generate a single structure per input. Notably, the inference configuration for HF3 w/ AFsample adopts a more sophisticated multi-setting approach, according to the AlphaFold settings used in AFsample [12]. The complete inference specifications for HF3 with AFsample integrate three distinct hyperparameter settings as detailed in Table 1.
Settings Templates Dropout Recycles Ratio %
setting-1 Yes Yes 3 30
setting-2 No Yes 3 30
setting-3 No Yes 9 40
Table 1: Inference configurations of HF3 w/ AFsample. The term Templates indicates whether structural templates were employed. Dropout denotes whether the dropout mechanism was activated. Recycles signifies the number of recycling operations utilized, with a default value of 3. Ratio represents the proportion that this particular setting occupies within the entire sampling process.
Evaluation Data
Evaluation sets for protein–protein, protein–ligand, protein–RNA, and protein–DNA interfaces were constructed from all PDB entries released between May 1, 2022 and December 31, 2024, with each structure expanded to Biological Assembly 1. Interfaces were defined as pairs of entities with a minimum heavy-atom distance below 5 Å. Protein–antibody complexes were sourced from SAbDab [32] within the same date range, using symmetric units instead of Biological Assembly 1.
For targets collected from the PDB, complexes with resolution worse than 4.5 Å or exceeding 1400 tokens under our tokenization scheme were removed. Polymer–polymer interfaces were excluded if both polymers shared more than 40% sequence identity with two chains from the same PDB entry in the training set. For protein–ligand interfaces, the following criteria were applied: (1) only ligands with CCD codes absent from the training set were retained; (2) covalently bound ligands, including those involved in glycosylation, were excluded; (3) ligands containing five or fewer atoms or occurring in ten or more PDB entries were removed; (4) only ligands with molecular weights between 100 and 900 Da were retained; (5) ligands were required to exhibit a ranking_model_fit score of at least 0.5, as reported in the RCSB structure validation dataset, indicating above-median model quality for X-ray crystallographic structures[33]; and (6) binding pockets were required to include between 5 and 100 protein residues within 5 Å of the ligand.
We clustered the remaining targets by grouping proteins with nine or more residues at 40% sequence identity, while nucleic acids and proteins with nine or fewer residues were clustered at 100%, using MMseqs2 with a minimum coverage of 80% and default clustering mode. Each interface was assigned a binary, order-independent cluster ID based on entity pairs—(polymer1_cluster, polymer2_cluster) for polymer–polymer interfaces and (polymer_cluster, ligand_CCD-code) for protein–ligand interfaces. Evaluation was performed on one representative entry per cluster.
Protein–antibody complexes were sourced from SAbDab [32], including only those with resolution better than 9 Å, containing antigen chains, and with antigen sequence identity less than 40% to any chain in the training set.
Evaluation Metrics
To evaluate structure prediction performance across different interaction types, we adopt distinct metrics tailored to the characteristics of each molecular interface.
Protein–protein complexes, including protein–antibody interactions, are evaluated using DockQ[34], which integrates interface RMSD, FNAT, and FNAS to provide a reliable summary of interface quality. For protein–antibody complexes specifically, all antibody chains are treated collectively as the “ligand”, and DockQ is computed over the interface between the antibody and the rest of the complex using the DockQ v1 implementation.
Nucleic acid–protein interfaces, including both protein–RNA and protein–DNA complexes, are assessed using interface LDDT (iLDDT) [35], computed over atom pairs across different chains within a 30Å inclusion radius to accommodate the larger and more diffuse interaction footprints characteristic of nucleic acids.
Protein–ligand complexes are evaluated using pocket-aligned RMSD, which measures ligand pose accuracy after aligning the predicted structure to the binding pocket of the ground truth. The pocket is defined as all heavy atoms within 10Å of any heavy atom of the ligand in the ground truth structure, restricted to the primary protein chain—identified as the chain containing the most atoms within this radius. The $C^{\alpha}$ atoms of this pocket are used to perform a least-squares alignment between predicted and reference structures. After alignment, a symmetry-corrected ligand RMSD is computed over all heavy atoms of the ligand using RDKit’s Chem.rdMolAlign.CalcRMS[36], which aligns the ligands while accounting for molecular symmetry before computing the final deviation.