### v2

# 1]PaddleHelix Team, Baidu Inc

Abstract
The abstract serves both as a general introduction to the topic and as a brief, non-technical summary of the main results and their implications. Authors are advised to check the author instructions for the journal they are submitting to for word limits and if structural elements like subheadings, citations, or equations are permitted.
Abstract
Generating large ensembles of candidate conformations is standard for improving biomolecular structure prediction. Yet aimless sampling is inefficient and costly, producing many redundant conformations with limited diversity, so additional computation often yields little improvement. Here, we present HelixFold-S1, a guided planning approach that strategically targets the most informative regions of conformational space to produce accurate conformations. For each biomolecule, predicted inter-chain contact probabilities serve as a blueprint of the conformational space, guiding computational effort toward higher-probability, low-redundancy contacts that constrain structure generation. Across diverse biomolecular benchmarks, HelixFold-S1 achieves markedly higher structural accuracy than traditional unguided methods while reducing sampling requirements by an order of magnitude. Predicted contact probabilities also provide a rough indicator of prediction difficulty and sampling utility. These results demonstrate that guided planning reshapes conformational exploration and enables more efficient and accurate structural inference.
keywords:
Biomolecular structure prediction, Conformation space exploration, Planning-driven, Computational efficiency, Resource-aware modeling
Introduction
Biomolecular structure prediction lies at the heart of computational biology, enabling progress in drug discovery, protein engineering, and the study of molecular interactions. Deep learning has transformed this field, with recent breakthroughs [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11] exemplified by the AlphaFold [1, 2, 3] and RoseTTAFold series [4, 5]. Despite these advances, accurate prediction of biomolecular complexes remains challenging.
Improving the accuracy of a single prediction has proved difficult, and generating large ensembles of candidate conformations has therefore emerged as a simple yet powerful way to boost performance. The AlphaFold series improves precision by model ensembling by combining predictions from five or twenty-five networks, and AlphaFold3 [3] further shows that extensive sampling can markedly improve protein–antibody interface modelling. Recent studies [12, 13, 14, 15] have explored large-scale sampling, generating thousands of candidate structures to expand conformational diversity. Approaches such as AFSample [12] and AFSample2 [13] introduce stochasticity during inference, using dropout and random template masking to generate alternative conformations, delivering notable gains in predictive accuracy. Other strategies [16, 17, 18, 19] diversify predictions by repeatedly sub-sampling the input multiple sequence alignments (MSAs), offering complementary routes to probe alternative structural states.
Figure 1: Overall framework of HelixFold-S1. a, Aimless exploration in conventional conformation sampling, often concentrated within limited regions of conformational space. b, Guided exploration in planning-based sampling, directing sampling toward more informative conformations. c, d, Two-stage inference of HelixFold-S1: c, stage 1 predicts a blueprint of the conformational space; d, stage 2 generates conformations based on this blueprint. e, Network architecture of HF-S1, which incorporates a contact conditioning module and a contact prediction module into the HelixFold3’s model architecture. During training, the two modules are activated in a mutually exclusive manner.
Even though generating large ensembles of candidate structures can improve predictive accuracy, it comes at substantial computational cost. For instance, performing a thousand samplings may require an entire GPU day. The main reason such massive sampling is needed is that current sampling methods are largely unguided, leading to aimless exploration of conformational space. As a result, the produced conformations often cluster around a single peak of the conformational landscape (Fig. 1a), yielding only a limited set of highly similar structures. Consequently, most sampled conformations provide little additional information, and overall sampling efficiency remains low. Moreover, it is often unclear whether additional sampling will meaningfully improve predictions, since simpler targets benefit less from extensive ensembling, rendering much of the computation redundant. These challenges highlight the need for planning-driven strategies that can first illustrate the most potentially informative regions, and then intelligently navigate the conformational space to focus computational effort where it matters most, thereby achieving higher accuracy at far lower computational cost.
Here, we introduce HelixFold-S1, a guided sampling strategy that explores conformational space with exceptional efficiency and precision. Instead of relying on random, unguided sampling, HelixFold-S1 strategically targets the most informative regions of the conformational space, with the aim of sampling across multiple higher-probability regions (Fig. 1b). For each biomolecule, we first predict inter-chain inter-residue contact probabilities based on pairwise representations. Interpreting these distributions as a coarse-grained structural blueprint of the conformational space provides a reduced representation that highlights where meaningful interaction patterns are most likely to occur (Fig. 1c). This structural map then guides the sampling, with computational effort deliberately concentrated in regions associated with higher-probability, low-redundancy contacts that act as spatial constraints during structure generation. The resulting conformations are evaluated and ranked by their confidence scores, enabling both accurate and resource-efficient modeling of complex biomolecular assemblies (Fig. 1d).
We evaluated HelixFold-S1 across diverse biomolecular interfaces, including protein–protein, protein–antibody, protein–ligand, protein–RNA, and protein–DNA. Compared with conventional unguided sampling, HelixFold-S1 substantially improves prediction accuracy at equivalent sampling effort, with the largest gains observed for challenging protein–antibody complexes. Remarkably, it achieves comparable accuracy to traditional methods using an order of magnitude fewer sampling steps. The approach is generalizable to other folding models to enhance structural precision. Predicted inter-chain contact probabilities correlate strongly with structural difficulty and provide a practical guide for allocating sampling effort across targets. Guided sampling also improves exploration of the conformational space, producing more diverse and higher-quality ensembles. HelixFold-S1 is deployed on the PaddleHelix platform [^1] https://paddlehelix.baidu.com/app/all/helixfold3/forecast and is available for online use.
Results
Guided Sampling Strategy and Architecture of HelixFold-S1
The inference pipeline of HF-S1 comprises two stages. In the first stage, HF-S1 predicts a coarse blueprint of the conformational space (Fig. 1c) to identify the most critical inter-chain interaction sites. To this end, an additional Contact Prediction Module is employed to estimate inter-chain inter-token contact probabilities based on the pair representation, representing the likelihood that any two tokens are in spatial proximity, i.e., any atom pair lies within 5Å. This contact probability map provides a reduced, coarse-grained representation of the high-dimensional conformational space, serving as a structural blueprint to guide subsequent conformation sampling. In the second stage, conformations are generated guided by the predicted blueprint. Contacts are prioritized according to their predicted probabilities and sequentially selected from the blueprint. Each selected contact is introduced as an additional constraint via the Contact Conditioning Module, which processes these constraints and generates structures that attempt to satisfy them. By focusing on contacts with higher predicted probabilities, the model is more likely to produce multiple high-quality conformations that capture the most informative interactions. To reduce sampling redundancy, we employ a strategy called redundancy contact pruning (RCP). Under this approach, once a conformation is generated using a particular contact, any contacts that are already satisfied within the resulting structure are excluded from further selection. All generated conformations are subsequently ranked according to model confidence scores, with the top-ranked prediction designated as the final output.
The architecture of HelixFold-S1 (HF-S1) builds upon that of HelixFold3 (HF3) [9], which reproduces the biomolecular structure prediction capabilities of AlphaFold3 (AF3) [3], and further extends its capacity to more intelligently explore conformational space through the addition of two contact-centric modules: the Contact Prediction Module (CPM) and the Contact Conditioning Module (CCM), as depicted in Fig. 1e. A contact is considered to exist between two tokens if any pair of atoms from the corresponding tokens lies within 5Å. The Contact Prediction Module computes a contact probability matrix across all token pairs using the pair representations produced by the Pairformer, a specialized attention-based module designed to capture long-range pairwise token dependencies. The Contact Conditioning Module, incorporated into the Input Embedder, enables the model to learn contact constraints and generate structures conditioned on these constraints. During training, these two modules are activated in a mutually exclusive manner, forming a multi-task framework that alternates between contact probability estimation and contact-conditioned structure generation. This design allows the model to jointly learn inter-chain interactions and effectively leverage them during inference. During inference, the two modules are employed sequentially: in the first stage (Fig. 1c), the Contact Prediction Module estimates the overall contact landscape, and in the second stage (Fig. 1d), the Contact Conditioning Module generates conformations guided by the predicted contacts.
Improved Structural Accuracy across Complex Types
Figure 2: Structural performance and module evaluation of HelixFold-S1 across multiple complex types. a, Top-5 structural precision among 25 sampled conformations for various folding models. Benchmark complexes were collected from the RCSB PDB between January 1, 2022, and December 31, 2024, including protein–antibody (n=221), protein–protein (n=198), protein–ligand (n=238), protein–RNA (n=177), and protein–DNA (n=254) complexes. b, Relative improvements of HelixFold-S1 over HelixFold3 across different complex types in Top-1, Top-5, and Top-25 precision. c, Comparison of HF-S1 with various baseline models on protein–antibody complexes released in 2024 (n=74), showing that extensive sampling progressively improves structural accuracy. Some methods were sampled 1000 times, while others used 100 samples. d, Performance comparison between HF3 w/ AFSample and HF-S1 across different antibody types on the same 2024 protein–antibody complexes (n=74), using 1000 samples. e, Performance of the contact prediction module in HF-S1, evaluated using the area under the precision–recall curve (AUPRC). f, Effectiveness of the contact conditioning module, measured as the fraction of predicted structures satisfying contact constraints. g, Example structures predicted by HelixFold-S1 using two contact constraints (PDB ID: 8ozb).
To systematically evaluate the performance of HF-S1 across a range of biologically relevant complex types, we constructed a test set comprising protein–antibody (n=221), protein–protein (n=198), protein–ligand (n=238), protein–RNA (n=177), and protein–DNA (n=254) complexes, collected between 2022.01.01 and 2024.12.31 from the RCSB PDB [20]. To minimize overlap with the training data, all test samples were selected to have low sequence identity to the training set. Sequences were clustered by similarity [21], and one representative per cluster was randomly chosen to ensure diversity and reduce redundancy. For protein–ligand complexes, any ligands that appeared in the training data were further excluded.
We compared AF3-like folding models [^2] AlphaFold3 is not available for commercial use and thus could not be tested. In addition, part of our benchmark overlaps with the Boltz-2 [22] training set, and Boltz-2 was therefore excluded., namely HF3 [9], Protenix [23], and Chai-1 [24], under different sampling strategies. For HF3, we examined two sampling-augmented variants: one following the AFSample strategy [12] (HF3 w/ AFSample) and another implementing our guided planning approach (HF-S1). We further transferred HF-S1–predicted contacts into Protenix and Chai-1, yielding Protenix w/ S1 and Chai-1 w/ S1, to assess their effect on alternative folding models. To enable parallel evaluation, Protenix w/ S1 and Chai-1 w/ S1 were run without redundancy contact pruning (RCP), which requires sequential sampling, whereas HF-S1 fully applies RCP during sampling. Because HF-S1 may employ a distinct definition of contacts from those folding models, the baseline models may not reach their best performance under this evaluation. Moreover, Chai-1 currently supports contact constraints only for protein–protein complexes, and results are therefore reported for protein–antibody and protein–protein systems only. For all methods, we generated 25 predicted structures per target and ranked them using each method’s confidence score.
The precision of the Top-5 ranked structures is shown in Fig. 2a. HF-S1 consistently outperforms both HF3 and HF3 w/ AFSample across all complex types, with the magnitude of improvement highest for protein–antibody complexes, followed by protein–protein and protein–RNA/DNA interfaces, and smallest for protein–ligand binding sites. The largest gains are observed for protein–antibody complexes, where HF-S1 achieves over a 55% improvement relative to HF3 and 33% relative to HF3 w/ AFSample. This likely reflects the structural diversity of antigen-binding sites, which makes conventional sampling less effective at capturing native-like geometries, whereas HF-S1’s guided sampling strategy increases the likelihood of generating correct interfacial contacts. For protein–protein and protein–RNA/DNA interfaces, HF-S1 provides moderate but clear improvements. By enforcing constraints on individual interfacial contacts, guided sampling helps capture correct local interactions, although single-contact constraints alone are not sufficient to fully determine the global geometry of these larger, more complex interfaces. Finally, protein–ligand binding sites are highly constrained, so while HF-S1 can still improve predictions, the relative gains are smaller compared with other complex types. Moreover, across different folding models, including HF3, Protenix, and Chai-1, incorporating the guided sampling strategy of HF-S1 leads to consistent performance gains, underscoring its broad effectiveness. In Fig. 2b, we further demonstrate that the performance advantage of HF-S1 over HF3 is consistently observed across different Top-K precisions of the sampled conformations. The improvements are pronounced for Top-1, Top-5, and Top-25, underscoring HF-S1’s robustness across multiple ranking thresholds, although the Top-1 gains exhibit greater variability, potentially reflecting sensitivity to the selection of conformations based on the confidence score.
Extensive sampling is particularly effective for protein–antibody interfaces [23], and HF-S1 shows the largest gains on these targets. To systematically evaluate the benefit of extensive sampling on these interfaces, we curated a dataset of 74 complexes released after 2024 and with residues less than 800 for efficiency, ensuring no overlap with Boltz-2’s training data [22]. Sampling was conducted with top-performing methods evaluated 1,000 times per complex, while others were sampled 100 times per complex. As shown in Fig. 2c, Top-1 precision on protein–antibody interfaces generally improves with increased sample size across all methods, particularly for those incorporating advanced strategies such as HF-S1 or AFsample. Importantly, HF-S1 achieves comparable precision to HF3 with AFsample using only 10 samples. This indicates that 10 HF-S1 samples are sufficient to reach the precision level that HF3 with AFsample requires 1,000 samples to achieve, representing just 1% of the computational cost. To further assess performance across different antibody types, Fig. 2d presents Top-1 precision for HF-S1 and HF3 with AFsample across canonical antibodies, nanobodies, and scFv based on the full 1,000-sample evaluation. HF-S1 demonstrates consistent improvements in Top-1 precision over HF3 w/ AFsample across all categories.
The Contact Prediction Module (CPM) and Contact Conditioning Module (CCM) are two key additions in HF-S1 built upon the HF3 architecture, and the accuracy of HF-S1 depends on both. To evaluate the CPM, we assessed the accuracy of its predicted contact probability matrices. For each target, token pairs corresponding to true contacts in experimental structures were treated as positives, and all others as negatives. We then computed the area under the precision–recall curve (AUPRC) between predicted and ground-truth contact maps, averaging across all targets. As a baseline, we derived a posterior contact probability matrix for HF3 by aggregating distances from sampled conformations, with each entry defined as the inverse of the minimum inter-residue distance. Compared with HF3, HF-S1 consistently produced more accurate contact probabilities across diverse molecular types (Fig. 2e). Protein–antibody complexes showed the lowest AUPRC, reflecting the difficulty of this prediction scenario, whereas protein–ligand complexes achieved the highest AUPRC, indicating a relatively simpler task. This observation is consistent with the results in Fig. 2a, where protein–antibody interfaces benefit most from advanced sampling strategies, while protein–ligand interfaces show smaller gains under the same approach. To examine the CCM, we evaluated whether HF-S1–predicted conformations adhered to specified contact constraints. We defined the contact satisfaction rate as the fraction of predicted structures in which the given contacts were realized. Across most test cases, HF-S1 achieved satisfaction rates above 70% (Fig. 2f), demonstrating its ability to effectively incorporate contact priors into structure prediction.
Fig. 2g illustrates a representative case. Two inter-chain residue pairs (R43–N26 and V103–L28) were selected from HF-S1’s predicted contact map and applied as spatial restraints during folding, yielding two distinct binding conformations that both satisfied the defined contact distances (R43–N26 $\approx$ 3.1 Å; V103–L28 $\approx$ 4.2 Å).
Contact Probability as an Indicator of Prediction Difficulty and Sampling Utility
Figure 3: Contact probability as an indicator of prediction difficulty and sampling utility. a, Correlation between target-level contact probability (maximum value in the predicted matrix) and structural precision across interface types, evaluated using the Top-5 of 25 sampled conformations. b, Distribution of predicted contact probabilities across interface types, constructed from all benchmark targets in each category. c, Structural improvement from single to extensive sampling by HF-S1 across contact probability groups, measured using the best-performing conformations, showing larger gains for lower-probability interfaces. d, Relative Top-5 precision (of 25 sampled conformations) of HF-S1 compared with HF3 across contact probability levels, indicating stronger benefits on lower-probability targets. e, Predicted contact probabilities reflect residue physicochemical preferences.
The contact probability matrix predicted by HF-S1 serves a dual role: it not only informs the structural sampling strategy but also provides insight into the intrinsic difficulty of the structure prediction task, as well as the potential benefits of multiple sampling.
We first show that predicted contact probabilities can serve as a proxy for estimating target difficulty. Specifically, we defined the target-level contact probability as the maximum value in the predicted contact probability matrix for each target and analyzed its relationship with the Top-5 precision of HF-S1 using 25 samples (Fig. 3a). Targets were stratified into high, medium, and low groups, revealing a strong correlation across datasets: lower-probability targets consistently yielded poorer predictions, while higher-probability targets were predicted with high precision. This indicates that low contact probability generally corresponds to more difficult targets, which often require additional sampling to achieve accurate predictions. When grouped by complex type (Fig. 3b), a similar trend is observed: Protein–protein targets typically show higher target-level contact probabilities, while protein–antibody complexes cluster in the lower range, indicating greater structural uncertainty. Protein–ligand complexes consistently exhibit high contact probabilities, suggesting that HF-S1 can often localize ligand binding sites with high confidence. In contrast, protein–RNA and protein–DNA complexes display a broader distribution, reflecting greater variability in prediction difficulty across these categories.
We next directly examined whether target-level contact probabilities are predictive of the benefits of multiple sampling, or sampling utility, which is valuable for efficiently allocating computational resources. Targets were grouped based on their target-level predicted contact probability, and we analyzed the precision improvements of the best-performing multi-sample prediction relative to single-sample predictions (Fig. 3c). We found that targets with intermediate contact probabilities achieved the greatest improvements, while those with low or high probabilities saw more modest gains. Notably, although lower-probability targets did benefit from sampling, their improvements were generally smaller than those in the intermediate group. This trend is intuitive: for targets with high contact probabilities, accurate structures can often be recovered from a single sample, leaving limited room for further enhancement. For lower-probability targets, the predicted contact maps are weak across the board, suggesting that a much larger number of samples may be needed to identify accurate structures. In contrast, intermediate cases offer partial yet informative contact signals, enabling the model to better explore the structural landscape and refine its predictions through sampling. As protein–ligand complexes consistently exhibit high target-level contact probabilities, they did not follow this trend.
We further examined performance across different contact probability groups by comparing HF-S1 (guided sampling) with HF3 (aimless sampling) (Fig. 3d). HF-S1 achieves the largest gains in the lower-probability group, while improvements are smaller in the higher-probability group. This trend suggests that targets with low contact probabilities are generally more challenging to model, and thus benefit more from guided exploration. In contrast, high-probability targets are easier to fold accurately, leaving less room for improvement.
Finally, we analyzed predicted contact probabilities across residue–residue pairs in the protein–protein dataset. Left panel of Fig. 3e: the averaged contact probability matrix across all residue pairs, with the pairs showing the highest probabilities indicated and those with the lowest probabilities highlighted. Right panel of Fig. 3e: compositional differences between the top 10 and bottom 10 pairs ranked by mean probability. Hydrophobic pairs such as Leu–Leu and Met-Met tend to show higher predicted probabilities, which could reflect their relatively frequent occurrence in tightly packed, energetically favorable environments. Conversely, polar or charged residues, such as Asp, Glu, and Lys, often exhibit lower probabilities, perhaps due to their general solvent exposure or context‐specific interaction tendencies. These patterns indicate that contact prediction models may capture fundamental physicochemical preferences underlying residue interactions [25, 26, 27, 28].
Guided Sampling Improves Exploration of Conformation Space
Figure 4: Guided sampling improves exploration of conformational space. a, Target-level standard deviation of structural precision among sampled conformations for HF3 and HF-S1. Each point represents a target; higher values indicate greater conformational diversity. b, Fraction of targets with higher standard deviation for HF-S1, higher for HF3, or comparable, summarized across different interface types.c, Structural precision as a function of the number of sampled conformations. d, Distribution of ranking confidence and structural precision of sampled structures for two representative complexes (PDB: 8PE1 and 8HHM). HF-S1 shows a broader distribution of structural precision, whereas HF3 predictions are more concentrated. Right panels show the corresponding predicted structures by HF-S1 and HF3 compared to experimental structures.
We hypothesize that guided sampling improves the exploration of conformational space, increasing the likelihood of near-native structures. To test this, we first compared HF-S1 (guided planning) with HF3 (aimless sampling) by calculating the standard deviation (std) of precision scores across sampled conformations for each target, where higher values indicate greater structural diversity (Fig. 4a). HF-S1 generally shows higher std values, with data concentrated in the upper-left region of the plot, reflecting more diverse ensembles. Targets were then grouped based on the difference in std between HF-S1 and HF3 (Fig. 4b): for protein–ligand complexes, a difference $\geq 1$ indicates HF-S1 std higher, $\leq-1$ indicates HF3 std higher, and intermediate values are comparable; for all other complex types, the threshold is 0.1. Across complex types, most targets fall into the HF-S1 std higher category, particularly protein–antibody, protein–RNA, and protein–DNA complexes, where the HF-S1 std higher category accounts for roughly twice as many targets as the HF3 std higher category. This enhanced diversity suggests that guided planning effectively directs sampling toward multiple plausible interaction geometries rather than converging prematurely on a single local mode.
To further examine how guided planning shapes the sampling trajectory, we tracked the average precision of all conformations accumulated up to each sampling step $K$ (Fig. 4c). For HF3, which performs aimless sampling, precision remains nearly constant across steps, as the order of sampled conformations is effectively random. In contrast, HF-S1 exhibits a gradual decline in precision as sampling proceeds, reflecting its design to explore progressively less probable contact configurations. The higher precision observed in the early sampling stages arises from the use of higher-probability contacts as structural constraints, which guide the generation of more accurate conformations. Notably, the first few conformations generated by HF-S1 already outperform those of HF3 across most interface types, except protein–DNA complexes. This behavior highlights that leveraging contact probabilities as structural constraints not only improves efficiency but also prioritizes structurally meaningful regions of the conformational landscape.
We further illustrate the differences between the aimless sampling strategy of HF3 and the guided sampling strategy of HF-S1 using two representative examples (Fig. 4d). The first example is the complex structure of fungal $\beta$-1,3-glucanosyltransferases (Gel4) and Nb4 nanobody, where the nanobody binds to a dissimilar CBM43 domain of Gel4 across fungal species (PDB ID: 8pe1) [29]. The second example is a protein-RNA heterodimer, illustrating the interaction between the Cas12m2 protein and crRNA (PDB ID: 8hhm) [30]. For each case, we examined the distribution of model-predicted confidence scores and interface-level precision metrics across the sampled conformations. HF-S1 consistently generates a broader spectrum of structures, spanning a wide range of confidence levels and precision values, including a notable fraction of high-accuracy predictions. In contrast, HF3 tends to sample conformations within a narrower confidence and precision range, indicating more limited structural diversity and fewer high-quality candidates.
Discussion
Extensive conformational sampling has become a standard strategy for improving structural prediction accuracy. However, unguided sampling remains computationally inefficient and offers limited insight into target difficulty or the approximate computational effort required. We propose that more intelligent exploration of the conformational space can enable more efficient allocation of computational resources. By constructing a blueprint of the conformational landscape, sampling can be directed toward regions that are more informative, enhancing both efficiency and prediction accuracy across diverse structure types. In addition, predicted contact probabilities can provide a rough indication of target difficulty, helping users to gauge the scale of sampling likely needed for reliable convergence.
Despite these advances, challenges remain in assessing structural confidence. Current confidence scores do not always reliably identify the most accurate structures, as high-confidence predictions may not correspond to true structural accuracy while low-confidence predictions can still include near-native conformations (Fig. S3). Improving these metrics through sophisticated scoring functions or ensemble-based calibration could facilitate more effective candidate selection and reduce the risk of overlooking high-quality conformations. Moreover, guided planning itself can exhibit diminishing returns, because greedily selecting contacts from highest to lowest probability may progressively sample less informative regions (Fig. 4c). Incorporating early-stopping strategies based on the saturation of higher-probability contacts or diminishing improvements in predicted structural quality could help maintain sampling efficiency. In addition, maintaining diversity among sampled conformations remains critical. Current approaches that sequentially exclude contacts from previously sampled structures reduce redundancy but require generating conformations one by one. Although this strategy improves performance (Fig. S4), further approaches are needed to anticipate which contacts to select in order to preserve diversity more efficiently.
Intelligently exploring the conformational space remains a critical and worthwhile challenge. More sophisticated search strategies, such as Monte Carlo tree search (MCTS) or other reinforcement learning-based planning methods, could be applied to systematically navigate this space. By dynamically balancing exploration and exploitation, these approaches may enable more efficient identification of high-quality conformations while preserving diversity, potentially reducing computational cost and enhancing the robustness of structure prediction across diverse biomolecular targets.
Methods
Model Architecture
HF-S1 builds upon the HF3 architecture and is designed to support two complementary tasks: inter-chain contact prediction and contact-conditioned structure prediction. To this end, HF-S1 introduces two additional components: the Contact Prediction Module and the Contact Conditioning Module, which extend the base architecture to enable contact-level reasoning and constraint-based structure generation, respectively.
The contact prediction task aims to estimate the inter-chain contact distribution of a given protein complex. Various input features—including sequence, multiple sequence alignment (MSA), and template information—are first encoded and then processed by a Pairformer module to generate single and pair representations. These pairwise representations are subsequently passed to a dedicated Contact Prediction Module, which includes a Pairformer Stack and performs a binary classification task on each element of the pairwise representation to output contact probabilities for each inter-chain token pair. A contact is defined as the presence of any atom pair between two tokens within 5Å in 3D space. The resulting pairwise contact probability matrix captures the inter-chain contact distribution of the complex and serves as an informative intermediate representation. By operating in the simplified space of contacts—rather than directly in the complex, high-dimensional structural space—the model achieves greater computational efficiency and facilitates the contact-conditioned structure prediction task.
The contact-conditioned structure prediction task introduces a Contact Conditioning Module to incorporate external contact constraints. These constraints are represented as a binary matrix $\{c_{ij}\}$, where $c_{ij}\in{0,1}$ indicates whether token pair $(i,j)$ is in contact (1 if any atom pair is within 5Å, and 0 otherwise). This matrix is projected through a linear layer and then fused into the pairwise activations within the Input Embedder of the model. During training, for each complex, 0–10 inter-chain contacts are randomly sampled from the contact set extracted from its ground-truth structure and provided as input. During inference, contact constraints are sampled from the contact probability matrix produced by the contact prediction task; the specific sampling strategy is described in a later section. The model learns to utilize the provided contact information to enhance structure prediction accuracy.
Training Regime
Parameters of the newly introduced Contact Prediction Module and Contact Conditioning Module are randomly initialized, while the remaining parts of the model inherit weights from the pretrained HF3. Fine-tuning is performed using the same training dataset as HelixFold3, which includes Protein Data Bank (PDB) [20] structures released before September 30, 2021, supplemented with self-distillation data to enhance generalization. The training follows a three-stage fine-tuning strategy: the first stage focuses on the contact-conditioned structure prediction task to improve complex structure accuracy with inter-chain contact constraints; the second stage adds the contact prediction task, jointly optimizing the model for both tasks; the third stage follows the same setting of the second stage but extends to larger crop size.
In the first stage, the model is trained exclusively on the contact-conditioned structure prediction task. For each training sample, inter-chain contacts are extracted from experimentally determined complex structures to form a ground-truth contact set $\mathcal{C}$. This set contains all inter-chain token pairs where at least one atom from each token lies within 5Å in three-dimensional space. Contact conditioning is applied with 70% probability: 1–10 token pairs are uniformly sampled from $\mathcal{C}$ and provided as binary contact constraints. In the remaining 30% of samples, no contact constraints are used, which helps maintain the model’s ability to predict structures without external guidance.
In the second stage, the contact prediction task is introduced, and the model is fine-tuned on both tasks simultaneously. The contact prediction task aims to estimate the probability that each inter-chain token pair is in atomic contact, serving as a basis for generating contact constraints during inference. During this task, the Contact Conditioning Module is not activated, and the model relies solely on encoded sequence, MSA, and template features to infer contact patterns. To supervise this task, a binary classification loss is applied over all inter-chain token pairs:
$\mathcal{L}_{\text{contact}}=\frac{1}{|\mathcal{P}|}\sum_{(i,j)\in\mathcal{P}}\text{cross\_entropy}(p_{ij}^{contact},y_{ij}^{contact}).$
Here, $\mathcal{P}$ denotes the set of all token pairs $(i,j)$ such that token $i$ and token $j$ belong to different chains. $p_{ij}^{contact}$ is the predicted probability of contact between tokens $i$ and $j$. $y_{ij}^{contact}=1$ if $(i,j)\in\mathcal{C}$ (i.e., the token pair is in contact), and $y_{ij}^{contact}=0$ otherwise. During training, half of the samples are used for the contact-conditioned structure prediction task, following the protocol established in the first stage, while the other half are dedicated to training the contact prediction task, which guides the model to estimate inter-chain contact probability distributions.
The loss function largely follows the original AF3/HF3 formulation, with an additional contact loss term introduced during fine-tuning:
$\mathcal{L}_{\text{loss}}=\alpha_{\text{confidence}}\mathcal{L}_{\text{confidence}}+\alpha_{\text{diffusion}}\mathcal{L}_{\text{diffusion}}+\alpha_{\text{distogram}}\mathcal{L}_{\text{distogram}}+\alpha_{\text{contact}}\mathcal{L}_{\text{contact}},$
with hyperparameters $\alpha_{\text{confidence}}=0.01$, $\alpha_{\text{diffusion}}=4$, and $\alpha_{\text{distogram}}=0.3$. The contact loss coefficient $\alpha_{\text{contact}}$ is set to 1 during training samples used for the contact prediction task and 0 during samples used for the contact-conditioned structure prediction task. The definitions of all other loss terms remain consistent with those in AF3.
All stages use the Adam optimizer [31] with parameters $\beta_{1}=0.9$, $\beta_{2}=0.95$, and $\epsilon=10^{-8}$, and a learning rate of $2\times 10^{-4}$. The mini-batch size is fixed at 128 for all stages. The first fine-tuning stage consists of 10,000 training steps with a crop size of 384. The second fine-tuning stage extends to 20,000 steps, also with a crop size of 384. The third stage continues training for an additional 3,000 steps with an increased crop size of 640.
Inference Regime
The inference process of HF-S1 (illustrated in Fig. 1a) consists of three stages: Contact Prediction, Contact Sampling, and Contact-Guided Structure Prediction and Ranking.
In the Contact Prediction stage, the contact prediction task is executed five times to reduce prediction variance, producing five contact probability matrices. These matrices are averaged element-wise to generate the final contact probability matrix, where each element $p_{ij}^{contact}$ represents the predicted contact probability between tokens $i$ and $j$. Only inter-chain contact probabilities are retained, with intra-chain contacts set to zero. For protein–antibody complexes, contact sampling is performed exclusively between the antigen chain and each antibody chain (heavy and light), excluding contacts between heavy and light chains.
In the Contact Sampling stage, inter-chain contacts are selected sequentially in descending order according to their predicted contact probabilities. Each selected contact is used as a binary constraint in the subsequent structure prediction step to generate diverse candidate structures. To improve sampling efficiency and avoid redundancy, two strategies are adopted: redundant contact pruning and enriched sampling of previously identified contact sets. We denote the sets of contacts extracted from previously predicted structures as $C_{1},C_{2},\ldots$, where each $C_{k}$ corresponds to the contacts obtained from the $k$-th predicted structure, following the same ground-truth extraction method described earlier. During sampling, redundant contact pruning excludes any candidate contact that overlaps with contacts already present in the union of all previously sampled sets $\bigcup_{i=1}^{k-1}C_{i}$. Here, overlapping means the candidate contact appears in any previously extracted contact set. This ensures that each newly sampled contact introduces novel constraints and helps maintain diversity among sampled structures. As sampling progresses, the predicted contact probabilities of remaining candidates naturally decrease. When these probabilities fall below a threshold (set as $0.2\cdot\max_{i,j}p_{ij}^{contact}$), the benefit of exploring new low-confidence contacts diminishes. At this point, instead of sampling new contacts, the algorithm enriches sampling by iterating through the existing contact sets $C_{1},C_{2},\ldots$ in order. Contacts are drawn from these sets cyclically to further exploit high-confidence information until the total sampling budget $S$ is reached.
In the Contact-Guided Structure Prediction and Ranking stage, each sampled contact is treated as a binary constraint and passed into the contact-conditioned structure prediction task to generate a candidate structure. A confidence score, named ranking_confidence, is computed for each structure, and the final prediction is selected as the one with the highest confidence among all candidates. Drawing inspiration from the AF3, we define the confidence score as a weighted average of the pTM and ipTM scores, with an additional penalty term for structural clashes. The score is computed as follows:
$\text{ranking\_confidence}=0.2\cdot\text{pTM}+0.8\cdot\text{ipTM}-1.0\cdot\text{has\_clash},$
where pTM represents the predicted TM-Score for the full complex, indicating the confidence for overall structural accuracy. ipTM represents the interface predicted TM-Score for the full complex, focusing on the accuracy of interfacial interactions. has_clash is a binary term indicating the presence of obvious clashes between polymer chains in the predicted structure. Detailed definitions of pTM, ipTM, and has_clash can be found in the AF3 paper [3].
We adopt consistent inference settings across structure prediction tasks, including our method and the baselines Boltz‑2[22], Protenix[23], and Chai‑1[24]. Each prediction is refined using 10 recycling iterations and 200 diffusion steps, where the diffusion module is run once to generate a single structure per input. In the corresponding figures, lines represent these average outcomes, while shaded areas indicate the variability between the two runs. Notably, the inference configuration for HF3 w/ AFsample adopts a more sophisticated multi-setting approach, according to the AlphaFold settings used in AFsample [12]. The complete inference specifications for HF3 with AFsample integrate three distinct hyperparameter settings as detailed in Table 1. Protenix w/ S1, Chai-1 w/ S1, and Boltz-2 w/ S1 follow the same workflow as HF-S1, differing only in the backbone structure prediction module employed during the final stage. Contact probability outputs from HF-S1 are ranked from highest to lowest, and the corresponding contact pairs are incorporated as external geometric constraints to guide structure generation. It should be noted that Chai-1 w/ S1 is restricted to protein–protein/antibody datasets, as it only accepts residue-residue contact constraints as input. All methods construct MSAs using their respective built-in sequence search tools.
Settings Templates Dropout Recycles Ratio %
setting-1 Yes Yes 3 30
setting-2 No Yes 3 30
setting-3 No Yes 9 40
Table 1: Inference configurations of HF3 w/ AFsample. The term Templates indicates whether structural templates were employed. Dropout denotes whether the dropout mechanism was activated. Recycles signifies the number of recycling operations utilized, with a default value of 3. Ratio represents the proportion that this particular setting occupies within the entire sampling process.
Evaluation Data
Evaluation sets for protein–protein, protein–ligand, protein–RNA, and protein–DNA interfaces were constructed from all PDB entries released between May 1, 2022 and December 31, 2024, with each structure expanded to Biological Assembly 1. Interfaces were defined as pairs of entities with a minimum heavy-atom distance below 5 Å. Protein–antibody complexes were sourced from SAbDab [32] within the same date range, using symmetric units instead of Biological Assembly 1.
For targets collected from the PDB, complexes with resolution worse than 4.5 Å or exceeding 1400 tokens under our tokenization scheme were removed. Polymer–polymer interfaces were excluded if both polymers shared more than 40% sequence identity with two chains from the same PDB entry in the training set. For protein–ligand interfaces, the following criteria were applied: (1) only ligands with CCD codes absent from the training set were retained; (2) covalently bound ligands, including those involved in glycosylation, were excluded; (3) ligands containing five or fewer atoms or occurring in ten or more PDB entries were removed; (4) only ligands with molecular weights between 100 and 900 Da were retained; (5) ligands were required to exhibit a ranking_model_fit score of at least 0.5, as reported in the RCSB structure validation dataset, indicating above-median model quality for X-ray crystallographic structures [33]; and (6) binding pockets were required to include between 5 and 100 protein residues within 5 Å of the ligand.
We clustered the remaining targets by grouping proteins with nine or more residues at 40% sequence identity, while nucleic acids and proteins with nine or fewer residues were clustered at 100%, using MMseqs2 with a minimum coverage of 80% and default clustering mode. Each interface was assigned a binary, order-independent cluster ID based on entity pairs—(polymer1_cluster, polymer2_cluster) for polymer–polymer interfaces and (polymer_cluster, ligand_CCD-code) for protein–ligand interfaces. Evaluation was performed on one representative entry per cluster.
Protein–antibody complexes were sourced from SAbDab [32], including only those with resolution better than 9 Å, and containing antigen chains. The antigen sequences were grouped into clusters based on a 40% sequence identity threshold. Subsequently, we retained only those clusters that contained no cases released prior to September 30, 2021. Further filtering was applied to select cases released within the period from May 1, 2022, to December 31, 2024. Ultimately, one case was chosen from each of the remaining clusters to constitute the evaluation set for protein-antibody analysis. For efficiency considerations in the extensive sampling test(as shown in Fig. 2c), we exclusively selected samples released between January 1, 2024, and December 31, 2024, and filtered out those with over 800 residues, ultimately retaining 74 samples.
Evaluation Metrics
To evaluate structure prediction performance across different interaction types, we adopt distinct metrics tailored to the characteristics of each molecular interface.
Protein–protein complexes, including protein–antibody interactions, are evaluated using DockQ [34], which integrates interface RMSD, FNAT, and FNAS to provide a reliable summary of interface quality. For protein–antibody complexes specifically, all antibody chains are treated collectively as the “ligand”, and DockQ is computed over the interface between the antibody and the rest of the complex using the DockQ v1 implementation.
Nucleic acid–protein interfaces, including both protein–RNA and protein–DNA complexes, are assessed using interface LDDT (iLDDT) [35], computed over atom pairs across different chains within a 30Å inclusion radius to accommodate the larger and more diffuse interaction footprints characteristic of nucleic acids.
Protein–ligand complexes are evaluated using pocket-aligned RMSD, which measures ligand pose accuracy after aligning the predicted structure to the binding pocket of the ground truth. The pocket is defined as all heavy atoms within 10Å of any heavy atom of the ligand in the ground truth structure, restricted to the primary protein chain—identified as the chain containing the most atoms within this radius. The $C^{\alpha}$ atoms of this pocket are used to perform a least-squares alignment between predicted and reference structures. After alignment, a symmetry-corrected ligand RMSD is computed over all heavy atoms of the ligand using RDKit’s Chem.rdMolAlign.CalcRMS [36], which aligns the ligands while accounting for molecular symmetry before computing the final deviation.
The precision of the Top-K ranked structures is defined as the highest accuracy achieved among the top K structures ordered by their confidence score.
Data Availability
To train HelixFold-S1, PDB can be downloaded at https://www.rcsb.org/docs/programmatic-access/file-download-services and AlphaFold Protein Structure Database as the distillation dataset can be downloaded at https://ftp.ebi.ac.uk/pub/databases/alphafold/v2/. The test set are filtered and clustered from PDB with conditions detailed in Methods. The protein-antibody complexes for test can be downloaded at https://opig.stats.ox.ac.uk/webapps/sabdab-sabpred/sabdab/.
Code Availability
The source code, trained weights, and inference scripts for HelixFold-S1 will be made publicly available upon acceptance. In addition, a web service for HelixFold-S1 is accessible at https://paddlehelix.baidu.com/app/all/helixfold3/forecast, providing efficient and accurate structure predictions.
Acknowledgments
This work was supported by the National Science and Technology Major Project (2023ZD0120803).
Supplementary Information
Appendix A Architecture of HelixFold-S1
HF-S1 is a modification of HF3 designed to support both the contact prediction task (Alg. 1) and the contact-conditioned structure prediction task (Alg. 2). HF-S1 extends HF3 by introducing two additional modules: the Contact Prediction Module (Alg. 3) and the Contact Conditioning Module, which integrates predicted contact information into the InputEmbedder (Alg. 4). All other modules remain consistent with those in AlphaFold3.
Algorithm 1 Main inference loop for contact prediction task
1:def ContactInferenceLoop($\{\mathbf{f}^{*}\}$, $N_{\text{cycle}}=4$, $c_{s}=384$, $c_{z}=128$) :
[…]
4:for all $c\in[1,\dots,N_{\text{cycle}}]$ do
5:  $\mathbf{z}_{ij}=\mathbf{z}_{ij}^{\text{init}}+\text{LinearNoBias}(\text{LayerNorm}(\hat{\mathbf{z}}_{ij}))$ $\triangleright$ $\mathbf{z}_{ij}\in\mathbb{R}^{c_{z}}$
6:  $\{\mathbf{z}_{ij}\}\mathrel{+}=\text{TemplateEmbedder}(\{\mathbf{f}^{*}\},\{\mathbf{z}_{ij}\})$
7:  $\{\mathbf{z}_{ij}\}\mathrel{+}=\text{MsaModule}(\{\mathbf{f}_{Si}^{\text{msa}}\},\{\mathbf{z}_{ij}\},\{\mathbf{s}_{i}^{\text{inputs}}\})$
8:  $\mathbf{s}_{i}=\mathbf{s}_{i}^{\text{init}}+\text{LinearNoBias}(\text{LayerNorm}(\hat{\mathbf{s}}_{i}))$ $\triangleright$ $\mathbf{s}_{i}\in\mathbb{R}^{c_{s}}$
9:  $\{\mathbf{s}_{i}\},\{\mathbf{z}_{ij}\}=\text{PairformerStack}(\{\mathbf{s}_{i}\},\{\mathbf{z}_{ij}\})$
10:  $\{\hat{\mathbf{s}}_{i}\},\{\hat{\mathbf{z}}_{ij}\}\leftarrow\{\mathbf{s}_{i}\},\{\mathbf{z}_{ij}\}$
11:end for
[…]
13:return $\{\text{p}_{ij}^{\text{contact}}\}$
Algorithm 2 Main inference loop for contact conditioned structure prediction task
1:def ContactConditionedInferenceLoop($\{\mathbf{f}^{*}\}$, $\{\text{p}_{ij}^{\text{contact}}\}$, $N_{\text{cycle}}=4$, $c_{s}=384$, $c_{z}=128$) :
[…]
5:for all $c\in[1,\dots,N_{\text{cycle}}]$ do
6:  $\mathbf{z}_{ij}=\mathbf{z}_{ij}^{\text{init}}+\text{LinearNoBias}(\text{LayerNorm}(\hat{\mathbf{z}}_{ij}))$ $\triangleright$ $\mathbf{z}_{ij}\in\mathbb{R}^{c_{z}}$
7:  $\{\mathbf{z}_{ij}\}\mathrel{+}=\text{TemplateEmbedder}(\{\mathbf{f}^{*}\},\{\mathbf{z}_{ij}\})$
8:  $\{\mathbf{z}_{ij}\}\mathrel{+}=\text{MsaModule}(\{\mathbf{f}_{Si}^{\text{msa}}\},\{\mathbf{z}_{ij}\},\{\mathbf{s}_{i}^{\text{inputs}}\})$
9:  $\mathbf{s}_{i}=\mathbf{s}_{i}^{\text{init}}+\text{LinearNoBias}(\text{LayerNorm}(\hat{\mathbf{s}}_{i}))$ $\triangleright$ $\mathbf{s}_{i}\in\mathbb{R}^{c_{s}}$
10:  $\{\mathbf{s}_{i}\},\{\mathbf{z}_{ij}\}=\text{PairformerStack}(\{\mathbf{s}_{i}\},\{\mathbf{z}_{ij}\})$
11:  $\{\hat{\mathbf{s}}_{i}\},\{\hat{\mathbf{z}}_{ij}\}\leftarrow\{\mathbf{s}_{i}\},\{\mathbf{z}_{ij}\}$
12:end for
[…]
16:return $\{\mathbf{x}_{i}^{\text{pred}}\},\{\text{plddt}_{l}^{\text{pred}}\},\{\text{p}_{\text{pae}}^{\text{pred}}\},\{\text{p}_{ij}^{\text{pred}}\},\{\mathbf{p}_{l}^{\text{resolved}}\},\{\text{p}_{ij}^{\text{distogram}}\}$
Algorithm 3 Contact prediction module
1:def ContactPredictionModule($\{\mathbf{s}_{i}\},\{\mathbf{z}_{ij}\}$) :
[…]
4:return $\{\text{p}_{ij}^{\text{contact}}\}$
Algorithm 4 InputEmbedder module with contact conditioning module
1:def InputEmbedder($\{\mathbf{f}^{*}\}$, $\{\text{c}_{ij}\}$, $c_{s}=384$, $c_{z}=128$) :
[…]
8:if $\{\text{c}_{ij}\}\neq\emptyset$ then
9:  $\mathbf{z}_{ij}^{\text{init}}\mathrel{+}=\text{LinearNoBias}(\text{c}_{ij})$ $\triangleright$ Contact conditioning module
10:end if
11:return $\{\mathbf{s}_{i}^{\text{init}}\}$, $\{\mathbf{z}_{ij}^{\text{init}}\}$
Figure S1: Performance of models on top-K metrics in 25 times sampling test. The top-K metric is defined as the highest metric value among the top-K predictions ranked by the confidence score. K values include 1, 5, 10 and 25.
Figure S2: Performance of models on success rates in the structure predictions. Three types of success rates, i.e., acceptable, medium and high, are calculated based on different thresholds of top-K metrics in 25 times sampling test. The thresholds are set as follows: 0.23, 0.49, and 0.8 for top-K DockQ scores of protein-antibody and protein-protein complexes; 0.25, 0.5, and 0.75 for top-K iLDDT scores of protein-RNA and protein-DNA complexes; and 2 Å, 1 Å, and 0.5 Å for top-K ligand RMSD values in protein-ligand complexes. K values include 1, 5, 10 and 25.
Appendix B Comparative Analysis of Top-K Metrics Across Different Methods
Fig. S1 presents a comprehensive visualization of the top-K metric results obtained from 25 sampling experiments for various methods across multiple datasets. The figure also highlights the statistical significance of the differences observed between the performance of different methods. From the results depicted in Fig. S1, it is evident that, under different K-value settings, methods incorporating S1 demonstrate a significant improvement in the top-K metrics compared to those without S1 integration. Specifically, on the Protein-DNA test set, while the top-1 iLDDT score for the model integrated with S1 does not show a statistically significant enhancement, notable improvements are observed in the top-5, top-10, and top-25 iLDDT metrics. This suggests that although the S1-integrated model is capable of identifying more accurate conformations, the precision limitations of the confidence score restrict the ability to fully reflect the effectiveness of its conformational search when solely comparing the highest-scoring conformation.
Fig. S2 provides a comparative analysis of the success rates achieved by different methods across various test sets. For each test set, we evaluated the success rates of each method at three distinct accuracy levels: acceptable, medium, and high. These accuracy levels correspond to different threshold values, the specifics of which are detailed in the figure caption. The analysis reveals that, compared to models without S1 integration, S1-integrated models generally exhibit a marked increase in overall success rates. Moreover, this improvement becomes more pronounced as the K-value increases. Additionally, the S1-integrated models predominantly enhance the success rates at the medium and high accuracy levels, indicating their capability to generate a greater number of medium- to high-quality conformations.
Appendix C Limitations of Ranking Scores
Figure S3: C-index scores assessing the ability of ranking confidence to correctly rank the quality of predicted structures generated by HF3 and HF-S1.
HF-S1 adopts the ranking confidence scoring strategy from AF3 to rank the diverse conformations generated during the ensemble process. To evaluate the effectiveness of this ranking method, we use the C-index metric to measure the correlation between the predicted ranking of conformations and their true quality ranking. For each test case, all conformations sampled by HF-S1 and HF3 are included to compute the corresponding C-index scores, which are then averaged across all cases. To avoid the influence of near-identical scores that may lead to ambiguous ordering, only conformation pairs with an absolute difference in true metric scores exceeding a threshold are considered. The threshold is 1 Å for ligand RMSD of protein-ligand complexes, and is 0.1 for DockQ or iLDDT of other complexes.
As shown in Fig. S3, the C-index scores of ranking confidence generally fall within the range of 0.5 to 0.7 across various test sets, where 0.5 corresponds to random ranking performance. This indicates a relatively weak correlation with the actual conformation quality. Although HF-S1 achieves slightly higher scores than HF3, the improvement is limited. Despite HF-S1’s ability to generate higher-quality conformations, the limited discriminatory power of ranking confidence makes it difficult to distinguish them effectively. This suggests that further refinement of the scoring strategy is critical to improving the selection of accurate conformations.
Appendix D Effect of Redundant Contact Pruning
Figure S4: Performance of HF-S1 with and without redundant contact pruning (RCP) on the protein-antibody 2024 dataset (n=103). For each case, 25 structures were sampled by each model. a, Accuracy comparisons among models. The max DockQ among the top-5 structures ranked by confidence score were used as the metric. b, The comparison of case-wise metric standard deviations (STD) between HF-S1 and HF-S1 without RCP.
HF-S1 employs a greedy sampling strategy based on descending contact probability and includes redundant contact pruning (RCP) to reduce sampling redundancy. We evaluated the impact of RCP on sampling diversity and model performance by conducting ablation experiments on protein-antibody 2024 dataset. Each model predicted 25 conformations for every case, and the highest DockQ among the top five conformations ranked based on the confidence score, was used to compare the performance differences between the models. As illustrated in Fig. S4a, the metrics of HF-S1 without RCP were lower than those of HF-S1 incorporating the RCP mechanism. Nevertheless, the results of HF-S1 without RCP still surpassed the levels of HF3 and HF3 with AFSample. This indicates that the RCP mechanism contributes to further amplifying the advantages of HF-S1. Fig. S4b compares the differences in the metric standard deviation (STD) between HF-S1 without RCP and HF-S1, aiming to demonstrate the impact of RCP on the diversity of model prediction results. Based on the relative differences in the STD of prediction scores between the two models across each case, the cases were categorized into three groups. Notably, the proportion of cases with a higher STD in HF-S1’s prediction metrics (43.7%) significantly exceeded that of the group with a higher STD in HF-S1 without RCP (11.7%). This suggests that HF-S1 facilitates an increase in the diversity of model prediction results, thereby enhancing the probability of the model achieving superior sampling.
Appendix E Additional Results on Protein–Antibody Structure Prediction
Epitope prediction plays a pivotal role in antibody therapeutic design. Building upon the observed improvements of HF-S1 in modeling protein–antibody complexes, we further evaluated its ability to identify epitope regions using protein–antibody 2024 dataset. Antibody Complementarity Determining Regions (CDRs) were annotated using ANARCI [37] with the Chothia numbering scheme. In this analysis, epitopes were defined as antigen residues that interact with at least one residue from any of the antibody’s CDRs, where an interaction is considered present if any pair of heavy atoms (one from each residue) is within 5Å. As shown in Fig. S5, HF-S1 consistently outperforms HF3 across all evaluation metrics, including F1 score, precision, and recall. Notably, HF-S1 demonstrates clear improvements, reflecting its enhanced ability to accurately identify epitope regions through contact-guided sampling.
Fig. S6 illustrates the growth trajectory of the oracle DockQ metric as a function of the number of sampling iterations on the antigen-antibody test set. The oracle DockQ represents the maximum DockQ value among all sampled conformations, providing an insight into the maximum achievable accuracy through extensive sampling by each model, independent of the precision limitations imposed by confidence scores. The results reveal a consistent upward trend in the oracle DockQ values for all models as the number of sampling iterations increases, aligning with the characteristics of a scaling law. Notably, the model integrated with S1 exhibits a significantly faster growth rate in its oracle DockQ metric compared to models without S1 integration. This accelerated progression underscores the capacity of the S1 model to amplify the benefits derived from increased sampling efforts. In essence, the S1-integrated model demonstrates a superior ability to leverage additional computational resources (i.e., more sampling iterations) to achieve greater gains in accuracy. The observed phenomenon highlights the potential of the S1 model to optimize the conformational search process, enabling it to explore and identify higher-quality conformations more efficiently.
Appendix F Assessment of Sequence Similarity Influence on HF-S1 Predictions
A common concern is whether target-level contact probability merely reflects the similarity between a test target and the training data—i.e., data overfitting—rather than capturing an intrinsic property of the target. To investigate this, we computed the sequence identity (defined in Supplementary Section L) between each test target and the training set, and examined its correlation with target-level contact probability. Fig. S7 show the relationship between sequence identity and the target-level contact probability predicted by HF-S1 across different datasets. Linear regression lines (red) with 95% confidence intervals (shaded areas) are displayed for each dataset. For protein–antibody, protein–protein, and protein–ligand, the correlations are negligible (Pearson $r=0.028$, 0.044-0.044, and $0.035$, respectively; all $p>0.05$), suggesting that HF-S1 predictions are largely insensitive to sequence similarity in these categories. While protein–DNA and protein–RNA show modest positive correlations ($r=0.230$ and $0.262$, respectively; $p<0.001$), indicating that higher sequence identity tends to be associated with slightly increased predicted contact probabilities in nucleic acid–binding systems. Overall, these results suggest that HF-S1 maintains robustness to sequence redundancy across protein- and ligand-related datasets, while moderate sequence-dependent effects may emerge in nucleic acid–interacting cases, possibly reflecting conserved recognition motifs or backbone–base contact patterns.
Appendix G Associations between residue types and the contact probability
A systematic survey of predicted residue–residue contact pairs across different confidence strata. Specifically, for each entry on protein–protein dataset, the top 50 and bottom 50 predicted contact pairs were extracted to represent the high- and low-probability regions, respectively. The frequencies of all residue–residue contact types were then aggregated across the dataset, and the top 10 most frequent contact pairs were selected for visualization. In high‐probability regions (Fig. S8 left), hydrophobic contacts dominate LEU–LEU pairs attain the highest scores, accounting about 19% of the high-confidence predictions, with additional hydrophobic combinations (e.g. ILE–LEU, LEU–PHE, LEU–VAL) collectively prevailing among the top predictions. Such concentration of hydrophobic interactions in the model’s strongest predictions aligns with the classical biophysical expectation that minimizing exposed hydrophobic surface area is a robust driver of protein protein interaction. In contrast, in low‐probability zones (Fig. S8 right), the model more frequently highlights charged residue pairs (e.g., GLU–LYS, ARG–GLU). These interactions, while less confidently predicted, may reflect the role of electrostatics and salt bridges in modulating binding specificity and affinity in context‐sensitive ways.
Figure S5: Capacity to identify epitopes.
Figure S6: Changes of orcale ensemble DockQs along the sampling numbers on protein-antibody 2024 dataset. For the efficiency of testing, only cases with total residue number less than 800 are included (n=74).
Figure S7: Correlations between sequence identity and contact probability predicted by HF-S1.
Figure S8: Associations between residue types and the contact probability predicted by HF-S1.
Appendix H Examples of Conformational Sampling Process
Fig. S9 provides illustrative examples of the conformational sampling process of HF-S1 across three types of complexes: protein-antibody, protein-protein, and protein-RNA. The prediction of each conformation involves a two-step procedure. Initially, a contact constraint is sampled from the contact probability map. Subsequently, this sampled constraint serves as a conditional input to predict a conformation that adheres to the specified constraint. The contact probability map exhibits a sparse characteristic, with the majority of regions displaying probabilities close to zero. This sparsity indicates that these regions are highly unlikely to form contacts. By leveraging the guidance provided by the contact probability map, our sampling approach effectively circumvents a substantial number of illegal contacts that would arise from random sampling, thereby ensuring a high quality of the sampled conformations. Our sampling method adopts a strategy that prioritizes contacts with higher probabilities and utilizes RCP mechanism to prevent redundant and overly concentrated sampling, thereby enhancing the diversity of the sampled contacts. A comparative analysis of the first five conformations sampled by HF-S1 for each case reveals significant differences among them. This variability can be attributed, to a certain extent, to the distinct guidance provided by different contact constraints. Each constraint steers the conformational search in a unique direction, leading to the exploration of diverse structural landscapes. The observed diversity in the sampled conformations underscores the effectiveness of our sampling strategy in exploring the conformational heterogeneity in biological complexes.
Figure S9: Examples of sampled contact constraints and their corresponding predicted structures from the first five samples of HF-S1. In each structure, distinct chains are color-coded for clarity. Below each model-generated structure, the positions of sampled contact constraints are annotated on the contact probability map using star symbols.
Appendix I Visualization of Predicted Structures by Different Models
Fig. S10 provides a detailed illustration of the changes in conformation prediction accuracy brought about by the application of the S1 mechanism. For each type of complex, a representative example is presented with corresponding PDB ID annotated on the figure, showcasing the highest-confidence-score conformation obtained from each model after 25 sampling iterations. In the figure, predicted conformations are depicted in cyan, while experimentally resolved conformations from the PDB are shown in gray. This color-coding scheme facilitates a clear visual distinction between the two, highlighting any discrepancies or similarities. In every example presented, the models with the S1 mechanism achieved significantly better metric scores compared to those without. This quantitative improvement is further corroborated by the visual alignment of the predicted and experimental conformations. Visually, the conformations predicted by models with the S1 mechanism exhibit a superior alignment with the experimentally resolved structures.
Figure S10: Examples of predicted structures generated by different models. For each case and model, the depicted structure is the highest-ranked (top-1) prediction, selected based on confidence scores from 25 model-generated structures.
Appendix J Analysis of Predicted Contact Probability Distributions
Fig. S11 presents the distribution of predicted contact probability values for the top 200 ranks across different types of biomolecular complexes. A notable observation is the significant variation in the highest contact probability values among different complex types. In protein-antibody complexes, the highest predicted contact probability is markedly lower compared to other complex types. Moreover, as the rank number increases, the probability values decline at a slower rate, resulting in a more long-tailed distribution. This pattern indicates a relatively high level of uncertainty in the model’s estimation of the docking interface positions. The long-tailed distribution suggests that there are numerous potential contacts with non-negligible probabilities, making it challenging to pinpoint the exact interaction sites.
The uncertainty observed in protein-antibody complexes can be attributed to the inherent characteristics of these molecules. Antigen-antibody interactions are known for their diversity, as antibodies can recognize and bind to a wide range of antigens with varying structures. Additionally, the lack of guidance from multiple sequence alignment (MSA) in these cases further exacerbates the uncertainty. MSA typically provides valuable information about conserved residues and potential interaction sites by comparing homologous sequences. However, in the absence of such alignment data, the model faces greater difficulty in accurately predicting contact probabilities. Traditional models, when faced with high levels of uncertainty such as those encountered in protein-antibody complexes, often provide a random estimate of the docking interface. In contrast, the S1 model, guided by contact probability values, conducts a more systematic exploration of the conformational space. By considering a range of possible contacts and their associated probabilities, the S1 model increases the likelihood of identifying high-quality conformations. This is particularly advantageous in cases where the interaction landscape is complex and uncertain, as it allows the model to capture the diversity of potential interactions and select the most probable ones.
Figure S11: The probability values of the top-200 predicted contacts for cases of different complex types.
Appendix K Comparative Distributions of Ranking Confidence and Structural Precision Between HF3 and HF-S1
Fig. S12 further illustrates case-wise distributions of metric values and confidence scores of more cases. For each biomolecular complex type studied, we selected three representative cases. In each case, both HF3 and HF-S1 performed 50 conformational samplings. Each scatter plot point represents a single sampling, with coordinates corresponding to its metric and confidence score. The results reveal that, across all complex types, HF-S1 generates metric values spanning a broader range than HF3, indicating greater conformational diversity. Additionally, HF-S1 produces a higher proportion of high-metric samplings, suggesting its superior ability to generate structures closely resembling experimentally determined conformations.
Another key observation is the positive correlation between high metric values and high confidence scores in many cases, enabling efficient filtering of high-quality samplings. However, there are many exceptions—such as in the protein-protein docking case 8k9e—where no clear relationship is observed. This discrepancy highlights the limitations of confidence scores in selecting high-accuracy structural predictions.
Figure S12: Case-wise distributions of metric values and confidence scores from HF3 and HF-S1 sampling. Each scatter plot displays the results of 50 structural samplings performed by one model on a single case. Here, each point represents the metric value and corresponding confidence score of an individual sampling.
Appendix L Definition and Formula for Sequence Identity
For a given complex $T$ consisting of polymer chains $\{C_{1},C_{2},\dots,C_{m}\}$, we define its sequence identity with respect to the training set $\mathcal{D}_{\text{train}}$ as
$\mathrm{ComplexId}(T,\mathcal{D}_{\text{train}})=\frac{1}{m}\sum_{i=1}^{m}\max_{S\in\mathcal{D}_{\text{train}}}\mathrm{SeqId}(C_{i},S),$
where the pairwise sequence identity between two sequences $A$ and $B$ is obtained via global Needleman–Wunsch alignment [38]:
$\mathrm{SeqId}(A,B)=\frac{M(A,B)}{L(A,B)},$
with $M(A,B)$ denoting the number of exact residue matches in the optimal alignment, and $L(A,B)$ the total alignment length (including gaps).